OpenStudy (anonymous):

What is LCM of p^2 and p^2+7p+10

OpenStudy (anonymous):

The last answer was wrong that I got

OpenStudy (anonymous):

p^2 + 7p + 10 = (p+5)(p+2) p^2 = p * p lcm = p

OpenStudy (anonymous):

eh, sorry